# Transformer Alignment in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. We regard LLMs as transforming embeddings via a discrete, coupled, nonlinear, dynamical system in high dimensions. This perspective motivates tracing the trajectories of individual tokens as they pass through transformer blocks, and linearizing the system along these trajectories through their Jacobian matrices. In our analysis of 38 openly available LLMs, we uncover the alignment of top left and right singular vectors of Residual Jacobians, as well as the emergence of linearity and layer-wise exponential growth. Notably, we discover that increased alignment positively correlates with model performance. Metrics evaluated post-training show significant improvement in comparison to measurements made with randomly initialized weights, highlighting the significant effects of training in transformers. These findings reveal a remarkable level of regularity that has previously been overlooked, reinforcing the dynamical interpretation and paving the way for deeper understanding and optimization of LLM architectures.
@@ 1 Introduction
Large language models (LLMs), as exemplified by BERT and GPT-series [Devlin et al., 2019, Brown et al., 2020], have revolutionized the field of natural language processing through their adoption of the transformer architecture [Vaswani et al., 2017]. Despite their widespread success, the internal mechanisms that underpin their performance are not fully understood.
Previous works viewed certain types of deep networks as implementing discrete, nonlinear dynamical systems, operating in high dimensions [Greff et al., 2016, Papyan et al., 2017, Ebski et al., 2018, Chen et al., 2018, Bai et al., 2019, Rothauge et al., 2019, Li and Papyan, 2023, Gai and Zhang, 2021, Haber and Ruthotto, 2017, Ee, 2017]. The term discrete reflects the network’s finite depth; nonlinear refers to the model’s nonlinear components; and dynamical is due to the residual connections spanning various layers. This motivates us to trace the dynamics of individual tokens as they traverse through the numerous transformer blocks, and to linearize the system through its Jacobian matrices.
1.1 Residual Alignment
Our investigation draws inspiration from a recent study by Li and Papyan [2023] on Residual Networks (ResNets) [He et al., 2016], which uncovered a phenomenon they termed Residual Alignment (RA), marked by several distinct characteristics in hidden representations of ResNets:
  • (RA1)
    Linear and equispaced trajectories in layer-wise progression of intermediate representations.
  • (RA2)
    Alignment of top left and right singular vectors in the linearizations of residual blocks.
  • (RA3)
    The rank of Residual Jacobians is at most the number of classes.
  • (RA4)
    Top singular values of Residual Jacobians scale inversely with depth.
@@ 1.2 Contributions
We uncover Transformer Alignment (TA), characterized by two properties (analogous to those of RA) that consistently emerge with training in 38 openly available base and fine-tuned LLMs (Appendix A.1) including the Falcon [Almazrouei et al., 2023], Llama-3 [AI@Meta, 2024], Llama-2 [Touvron et al., 2023], MPT [Team, 2023a, b], Mistral v0.1 [Jiang et al., 2023], Gemma [Team et al., 2024], NeoX [Black et al., 2022], Neo [Black et al., 2021, Gao et al., 2020], and Pythia [Biderman et al., 2023] families. The properties are given by
  • (TA1)
    Emergence of linearity and exponentially distanced trajectories of intermediate representations in layer-wise progression.
  • (TA2)
    Alignment of top left and right singular vectors in the linearizations of transformer blocks through their Residual Jacobians.
Due to the large vocabulary size relative to the embedding size in LLMs, Residual Jacobians maintain rank of at most the vocabulary size, and so the previous (RA3) immediately holds. Further, trained LLMs do not exhibit a consistent inverse scaling of top singular values with depth, and scaling varies considerably between models, demonstrating a lack of (RA4). In contrast to ResNets, the layer-wise progression of transformer representations is better described with an exponential evolution, rather than an equispaced linear progression.
Additionally, we explore the relationship between (TA2) and generalization, and uncover that increased alignment of Residual Jacobian singular vectors correlates with improved Open LLM leaderboard [Beeching et al., 2023] benchmark score (Figure 9), raising valuable insight to the mathematical properties of LLM transformers and their connections to generalization.
@@ 2 Background on Large Language Models
In the input layer, $l=0$, textual prompts undergo tokenization and are combined with positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token. When these embeddings are stacked together, they form a matrix:
$X^{0}=(x_{1}^{0},x_{2}^{0},\ldots,x_{n}^{0})\in\mathbb{R}^{d_{\text{model}}\times n}.$ (1)
@@
Here, $X^{l}=f_{\text{block}}^{l}(X^{l-1})$ denotes the embeddings after the $l^{\text{th}}$ block, consisting of causal multi-headed attention (MHA), a feed-forward network (FFN), and normalization layers (LN) with residual connections:
@@
$f_{\text{block}}^{l+1}(X^{l})$ $=X^{l}+h^{l+1}(X^{l})+\text{FFN}(g^{l+1}(X^{l})),$ (5)
where the MHA, LN, FFN are implicitly indexed by layer. In the final representation, an additional layer normalization is applied:
@@ 3 Methods
3.1 Suite of Large Language Models
Our study evaluates a total of 38 LLMs (24 base LLMs and 14 fine-tuned, see Appendix A.1) that were independently trained by various individuals and organizations. These models, provided through HuggingFace [Wolf et al., 2020], vary in terms of parameter budgets, number of layers, hidden dimensions, and training tokens. A summary of the models under consideration is presented in Table 1 of Appendix A.1. Measurements are made for models with varying parameter sizes with GPUs satisfying appropriate memory requirements, further detailed in Appendix A.3.
3.2 Prompt Data
We evaluate these LLMs using prompts of varying length, ambiguity, and context, sourced from the test set of ARC [Clark et al., 2018], GSM8K [Cobbe et al., 2021], HellaSwag [Zellers et al., 2019], MMLU [Hendrycks et al., 2021], Truthful QA [Lin et al., 2022], and Winogrande [Sakaguchi et al., 2019]. These datasets set the performance benchmarks on the HuggingFace Open LLM Leaderboard [Beeching et al., 2023] since they encompass a diverse set of language tasks. Post-prompting, we assess the models using several metrics, detailed in the following subsections.
3.3 Alignment of Singular Vectors of Residual Jacobians
To demonstrate the presence of (TA2), we examine the properties of the transformer blocks and the relationships between them. This is done by analyzing the linearizations of the blocks given by their Residual Jacobian matrices
$J_{l}^{i}=\frac{\partial}{\partial x_{i}^{l-1}}\left(h^{l}(X^{l-1})+\text{FFN\textsubscript{$l$}}(g^{l}(X^{l-1}))\right)_{i},$ (7)
@@ 3.3 Alignment of Singular Vectors of Residual Jacobians
defined for each $l=1,\ldots,L$, and $i=1,\ldots,n$. Note that this is the Jacobian matrix for each transformer block without the contribution from the skip connection from the input of the block, analogous to the quantities measured by Li and Papyan [2023].
@@
The singular value decompositions of these $J_{l}^{i}$ are computed, i.e., $J_{l}=U_{l}S_{l}V_{l}^{\top}$ (with superscript $i$, indicating the token, omitted for clarity), where $U_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ and $V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are the matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ is the singular value matrix.
@@
The matrices $A_{i,j,K}\coloneqq U_{j,K}^{\top}J_{i}V_{j,K}$ are plotted over all pairs of depths $i,j\in\{1,\ldots,L\}$, where $U_{j,K}$ and $V_{j,K}$ are the sub-matrices with columns that are the top-$K$ left and right singular vectors of $J_{j}$, respectively.
@@
To quantify the misalignment of singular vectors of the Residual Jacobians at depths $i$ and $j$, we measure how far the $A_{i,j,K}$ are from being diagonal. More precisely, we define the misalignment of $A_{i,j,K}$ as
$m(A_{i,j,K})=\lVert A_{i,j,K}-\text{Diag}(A_{i,j,K})\rVert_{F},$ (8)
@@
where $\text{Diag}(A_{i,j,K})$ is the matrix $A_{i,j,K}$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm.
3.4 Layer-Wise Exponential Growth
To demonstrate the presence of (TA1), we measure the expodistance of the hidden trajectories, defined below. In this context, each layer in the model corresponds to a unit time step of a dynamical system, so we denote $\dot{x}_{i}(t)=x_{i}(t+1)-x_{i}(t)$ the time derivative of the $i^{\text{th}}$ token. Assuming exponential growth of the embedding norms as they flow through the hidden layers, we can write,
$\|\dot{x}_{i}(t)\|$ $\approx e^{\alpha t}\|\dot{x}_{i}(0)\|$ (9)
for some fixed $\alpha\in\mathbb{R}$. We quantify the validity of the approximation in Equation (9) by measuring the coefficient of variation of $\alpha$ for each trajectory. Namely, following (9),
$\alpha_{t}^{i}\approx\frac{1}{t}\ln\left(\frac{\|\dot{x}_{i}(t)\|}{\|\dot{x}_{i}(0)\|}\right)$ (10)
for each time (layer) $t$ and token $i$. Under exponential growth, it is expected that $\alpha_{t}^{i}$ is independent of time. We then denote the expodistance (ED) of the trajectory of the $i^{th}$ token of a given sequence by
$ED_{i}=\frac{\text{Var}_{t}\alpha_{t}^{i}}{(\text{Avg}_{t}\alpha_{t}^{i})^{2}}$ (11)
This measurement is motivated by the parametrization discussed in Appendix A.2 and serves as a method to test the validity of the linearization presented in Equation (15).
@@ 3.5 Linearity of Trajectories
Linearity in intermediate embeddings (TA1) is quantified with the line-shape score (LSS), defined by Gai and Zhang [2021] as
@@
3.6 Visualization of Trajectories with PCA
@@ 4 Results
4.1 Alignment of Singular Vectors of Residual Jacobians
@@
In trained LLMs, we observe (TA2), that is, alignment of the top singular vectors of the Jacobians $J_{l}$ across depth (Figure LABEL:fig:paris_all_8to16 bottom row), evident in the distinct diagonal lines present in the matrix subplots. This phenomenon is consistently observed across LLMs, similar to (RA2), which was observed by Li and Papyan [2023] in ResNets. On the other hand, in untrained models (Figure LABEL:fig:paris_all_8to16 top row), there is no alignment between singular vectors of Residual Jacobians across different depths. There is alignment along the diagonal, however, because each Residual Jacobian is trivially diagonalized by its own singular vectors in $A_{i,j,K}$. This suggests that (TA2), the alignment of singular vectors of Residual Jacobians, emerges through training.

### v2

# Transformer Block Coupling and its Correlation with Generalization in LLMs

@@ Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. In this work, we trace the trajectories of individual tokens as they pass through transformer blocks, and linearize the system along these trajectories through their Jacobian matrices. By examining the relationships between these Jacobians, we uncover a transformer block coupling phenomenon in a variety of LLMs, characterized by the coupling of their top singular vectors across tokens and depth. Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters, namely parameter budget, model depth, and embedding dimension. We further investigate the emergence of these properties through training, noting the development of coupling, as well as an increase in linearity and layer-wise exponential growth in the token trajectories. These collective insights provide a novel perspective on the interactions between token embeddings, and prompt further approaches to study training and generalization in LLMs.
@@ 1 Introduction
In recent years, many openly available Large Language Models (LLMs) have been released, achieving state-of-the-art results on task-specific benchmarks. The abundance of models, each with differing architecture and training methodology, motivates comparing the underlying mechanisms that drive generalization.
Transformers (Vaswani et al., 2017) can be represented as discrete, nonlinear, coupled dynamical systems, operating in high dimensions (Greff et al., 2016; Papyan et al., 2017; Haber & Ruthotto, 2017; Ee, 2017; Ebski et al., 2018; Chen et al., 2018; Bai et al., 2019; Rothauge et al., 2019; Gai & Zhang, 2021; Li & Papyan, 2023). The term discrete reflects the network’s finite depth; nonlinear refers to the model’s nonlinear components; coupled refers to the interdependent token trajectories enabled by self-attention, and dynamical is due to the residual connections spanning various layers.
Throughout this work, we investigate whether we can identify structural characteristics that consistently occur in pretrained LLMs that are related to their generalization, as measured on the HuggingFace OpenLLM Leaderboard benchmark. To this end, we study and quantify the trajectories of token embeddings and their interactions, through linearizations of LLM layers given by the Jacobians of the associated transformer blocks.
This perspective raises several questions:
Q1. What regularity properties do these trajectories exhibit individually, and what are their relations with one another? More concretely, what is the relation between the Jacobians across different tokens and transformer layers?
Q2. How do the properties of hidden representations and their relations emerge with training?
Q3. Are any of these properties related to the generalization capabilities of LLMs?
Figure 1: Relationship between Coupling and Benchmark Score. The plot illustrates the relationship between average coupling (negative mis-coupling) and benchmark score across various base LLMs (that are not fine-tuned). The data shows a noticeable trend: increased alignment correlates with improved benchmark scores. The regression fit produces an $R^{2}$ value of $0.7214$.
@@ 1.1 Contributions
We investigate the motivating questions across several openly-available LLMs, most having over 1B parameters, trained by 6 independent organizations, with varying training methods and data (Appendix A.1). Through our experiments, we identify consistent patterns that define the transformer block coupling phenomenon.
1. Coupling. The singular vectors of the Jacobians of transformer blocks couple across depth (Figures LABEL:fig:paris_all_8to16, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix). Further, coupling across Jacobians emerges with training (Figures 2, LABEL:fig:paris_all_8to16, LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout).
2. Generalization. The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures 1, LABEL:fig:alignment_vary_k). Additionally, coupling is more strongly correlated with generalization than parameter budget, model depth, and token embedding dimension (Figure LABEL:fig:correlation_hyperparams).
3. Regularity. Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, 13, 15, 23, 24). Exponential growth occurs in contiguous token representations as a function of depth (Figures 14, 22, 23, LABEL:figs:expo-training).
We provide a new perspective on token embedding interactions within LLMs by examining transformers layers through their Jacobians. Our results display the effect of training on transformer blocks, and suggest potential approaches for promoting generalization in LLMs.
Figure 2: Emergence of Singular Vector Coupling with Training. Normalized coupling (Section 3.1) evaluated for $J_{i,j}^{d_{1}},J_{i,j}^{d_{2}}$, with $|d_{1}-d_{2}|\leq 5$. Measurements are taken for checkpoints of Pythia 12B, Pythia 6.9B, Pythia 1B (Biderman et al., 2023) at training steps $\{1,16,128,13\text{k},23\text{k},33\text{k},\ldots,143\text{k}\}$. Coupling is low at early training iterations, and increases with training, with Pythia 12B and 6.9B displaying greater coupling than Pythia 1B in final training steps.
@@ 2 Background on Large Language Models
We describe LLMs as a deep composition of functions that iteratively transform token embeddings. In the input layer, $l=0$, textual prompts undergo tokenization and are combined with positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token. When these embeddings are stacked, they form a matrix:
$X^{0}=(x_{1}^{0},x_{2}^{0},\ldots,x_{n}^{0})\in\mathbb{R}^{n\times d_{\text{model}}}.$ (1)
@@
$X^{l}=F_{\text{block}}^{l}(X^{l-1})$ denotes the embeddings after the $l^{\text{th}}$ block, consisting of causal multi-headed attention (MHA), a feed-forward network (FFN), and normalization layers (LN) with residual connections:
@@
$f^{l+1}(X^{l})$ $=h^{l+1}(X^{l})+\text{FFN}(g^{l+1}(X^{l}))$ (5)
$F_{\text{block}}^{l+1}(X^{l})$ $=X^{l}+f^{l+1}(X^{l}),$ (6)
where the MHA, LN, FFN are implicitly indexed by layer. Among many models (Appendix 1), an additional rotary positional embedding (RoPE, Su et al. (2023)) is applied in the MHA layer. In the final representation, typically an additional layer normalization is applied:
@@ 3 Methods
3.1 Coupling of Singular Vectors of Jacobians
Jacobians. Coupling is investigated through analyzing the linearizations of transformer blocks given by their Jacobians
$J_{t_{1}t_{2}}^{l}=\frac{\partial}{\partial x_{t_{1}}^{l-1}}\left(f^{l}(X^{l-1})\right)_{t_{2}},$ (8)
@@ 3.1 Coupling of Singular Vectors of Jacobians
defined for each layer $l\in\{1,\ldots,L\}$, and pair of tokens $t_{1},t_{2}\in\{1,\ldots,n\}$. Note that this is the Jacobian matrix for each transformer block without the contribution from the skip connection from the input of the block, similar to the quantity measured by Li & Papyan (2023), but extended for cross-token interactions. Also, due to the causal structure of the representations, $J_{t_{1}t_{2}}^{l}=0$ if $t_{1}>t_{2}$, so we restrict our attention to the case $t_{1}\leq t_{2}$.
@@
Singular value decomposition. The SVD of the Jacobians $J_{t_{1}t_{2}}^{l}$ are computed, i.e., $J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ (with superscripts $t_{1},t_{2}$, indicating the tokens, omitted for clarity), where $U_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ and $V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ is the singular value matrix.
@@
To measure coupling between Jacobians $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, define $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}\coloneq U_{l^{\prime},K}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime},K}$ for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$, where $U_{l^{\prime},K}$ and $V_{l^{\prime},K}$ are the sub-matrices with columns that are the top-$K$ left and right singular vectors of $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, respectively.
Figure 3: Transformer Block Coupling. Displayed is a visualization of the various types of transformer block coupling, along with brief instructions on computing both the Jacobians $J$ and coupling matrices $A$.
Depth-wise Coupling. To analyze coupling across transformer blocks, we fix $t$ (usually as the last token), and check alignment between $J_{tt}^{l}$ and $J_{tt}^{l^{\prime}}$ through $A_{t}^{ll^{\prime},K}$ (we write the single subscript $t$ for clarity) across layers $l,l^{\prime}\in\{1,\ldots,L\}$.
Token-wise Coupling. To study the coupling across tokens, we start by fixing two layers $l,l^{\prime}\in\{1,\ldots,L\}$, and analyze coupling across different cases of the $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}$. We start with the case where the input and output tokens are the same and compare $J_{tt}^{l}$ and $J_{t^{\prime}t^{\prime}}^{l^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$, which is the coupling across tokens for a token’s effect on its own trajectory. We also consider the context tokens’ impact on a trajectory by including coupling across cases with $t_{1}<t_{2}$. More specifically, we measure coupling between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}t_{2}^{\prime}}^{l^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$ (fixing the input token to be the same) and also between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}}^{l^{\prime}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$ (fixing the output token to be the same).
@@
Coupling measurement. To quantify the coupling of singular vectors of the Jacobians, we measure how far the $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ are from being diagonal. More precisely, we quantify the mis-coupling of $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ using $m:\mathbb{R}^{K\times K}\to\mathbb{R}$ defined by
$m(A)=\lVert A-\text{Diag}(A)\rVert_{F},$ (9)
@@
where $\text{Diag}(A)$ is the matrix $A$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm. For normalized comparison between models across training in Figure 2, we normalize $A$ by their norms
$\tilde{m}(A)=\frac{\lVert A-\text{Diag}(A)\rVert_{F}}{\lVert A\rVert_{F}}$ (10)
@@ 3.2 Linearity of Trajectories
Linearity in intermediate embeddings is quantified with the line-shape score (LSS), defined by Gai & Zhang (2021) as
@@
3.3 Layer-Wise Exponential Growth
We measure the presence of exponential spacing (expodistance) of the hidden trajectories. Assuming exponential growth of the embedding norms as they flow through the hidden layers, we can estimate,
$\|x_{i}^{l}\|$ $\approx e^{\alpha l}\|x_{i}^{0}\|=e^{\alpha}\|x_{i}^{l-1}\|$ (13)
for some fixed $\alpha\in\mathbb{R}$ over all layers $l=1,\ldots L$. We quantify the validity of the representation in Equation 13 by measuring the coefficient of variation of $\alpha_{i}^{l}$, given by
$\alpha_{i}^{l}\approx\ln\left(\frac{\|x_{i}^{l}\|}{\|x_{i}^{l-1}\|}\right),$ (14)
for each layer $l$ and token $i$. Under exponential growth, it is expected that $\alpha_{i}^{l}$ is independent of depth. We therefore denote the expodistance (ED) of the trajectory of the $i^{th}$ token of a given sequence by
$\text{ED}_{i}=\frac{\text{Var}_{l}\alpha_{i}^{l}}{(\text{Avg}_{l}\alpha_{i}^{l})^{2}}.$ (15)
This measurement is motivated by the parametrization discussed in Appendix A.4, as well as empirical evidence in Figure LABEL:figs:norm-expo-traininga, and serves as a method to test the validity of the linearization presented in Equation 17.
4 Evaluation
4.1 Suite of Large Language Models
Our study evaluates a total of 38 LLMs (24 base LLMs and 14 fine-tuned, see Appendix A.1) that were independently trained by various individuals and organizations. These models, provided through HuggingFace (Wolf et al., 2020), vary in terms of parameter budgets, number of layers, hidden dimensions, and training tokens. Moreover, we analyze the dynamics of each measurement throughout training by deploying the Pythia Scaling Suite (Biderman et al., 2023). A summary of the models under consideration is presented in Table 1 of Appendix A.1 and further details in Appendix A.5.
4.2 Prompt Data
We evaluate these LLMs using prompts of varying length, ambiguity, and context, sourced from the test set of ARC (Clark et al., 2018), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), Truthful QA (Lin et al., 2022), and Winogrande (Sakaguchi et al., 2019). This data sets the performance benchmarks on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) and provide a representative evaluation of performance on many language tasks.
@@ 5 Results
5.1 Coupling of Jacobians Across Depth
@@
In trained LLMs, we observe coupling of the top singular vectors of the Jacobians across depth (Figure LABEL:fig:paris_all_8to16 bottom row), evident in the low non-diagonal values with a visible diagonal present in the matrix subplots. This is consistently observed across various LLMs considered. On the other hand, in untrained models (Figure LABEL:fig:paris_all_8to16 top row), there is no coupling of Jacobians across different depths. There is coupling along the diagonal, however, because each Jacobian is trivially diagonalized by its own singular vectors. This, in addition to Figure 2, suggests that coupling across depth emerges through training.
5.2 Coupling of Jacobians Across Tokens
We also analyze the coupling of singular vectors of the Jacobians across tokens. First, we observe coupling between Jacobians where the input and output tokens are the same, that is, between $J_{l}^{tt}$ and $J_{l^{\prime}}^{t^{\prime}t^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$ (Figure LABEL:fig:coupling_inoutsame). This suggests that a token’s interaction on its own trajectory is coupled with that of other tokens. We also consider the context tokens’ interaction on a trajectory when fixing the input token $t_{1}$ (Figure LABEL:fig:coupling_fixedin), through $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$, and when fixing the output token $t_{2}$ (Figure LABEL:fig:coupling_fixedout), with $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$. Although generally present, the coupling in these two cases are less consistent, with variation in the strength of coupling between different token pairs. Again in untrained models (Figure LABEL:fig:coupling_inoutsame top row), there is a lack coupling of Jacobians across tokens.
5.3 Emergence of Coupling with Training
Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure LABEL:fig:paris_all_8to16 ) and across tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout). Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 12B, 6.9B, and 1B (Biderman et al., 2023) (Figure 2), and observe that coupling is generally low at initialization and increases significantly during early training steps.
5.4 Correlation with Generalization